OpenStudy (anonymous):

All of the following proportions are equivalent except a/b=c/d b/a=d/c a/d=c/b a/c=b/d

OpenStudy (turingtest):

we are missing a picture I presume

OpenStudy (anonymous):

nope

OpenStudy (turingtest):

we are missing something...

OpenStudy (nenadmatematika):

calistar something's missing in your problem for sure...:D

OpenStudy (anonymous):

srry thats the way the problem is, no pic no nothing ..thanxx anyway=)

OpenStudy (anonymous):

a/d=c/b(that's wrong)

OpenStudy (nenadmatematika):

aha I think I got it, one equality written in four ways, so which is false...if a/b=c/d than it's not possible that a/d=c/b

OpenStudy (turingtest):

I get it they each become ad=bc bc=ad ab=cd ad=bc so the third one